Enchanted Kingdom’s EKstreme Tower ride has been popular for daredevils. A group of senior high school students decided to take on the challenge and conquer the ride. After reaching the top the students brace their selves for the drop that will take for 3.2 seconds. Solve for the final velocity and the height that have reached before the drop happened


1
Expert's answer
2021-09-21T06:33:30-0400

Let's first find the height of the tower:


"h=\dfrac{1}{2}gt^2=\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(3.2\ s)^2=50.2\ m."

Finally, we can find the final velocity:


"PE=KE,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot50.2\ m}=31.4\ \dfrac{m}{s}."